Table represents Profit Contribution DEMAND Strong Weak $1,212 -$313 $801$422 Full Price Discount Suppose that management believes that the probability of strong demand is 0.70 and the probability of weak demand is 0.30. Use the expected value approach to determine an optimal decision. Also calculate the expected value of perfect information
Answer #1

1)here expected value of Full price =1212*0.7+(-313)*0.3=754.5

expected value of Discount =801*0.7+422*0.3=687.3

as expected value for Full price is highest ; therefore Full price is optimal decision.

2)

here expected value with perfect information=1212*0.7+422*0.3=975

threfore expected value of perfect information=expected value with perfect information-expected value without perfect information

=975-754.5=220.5